# The Signals of the Doomsday

### v3

@@ Abstract
The measured standard model parameters indicate that we might live in a false Higgs vacuum, though with a very long lifetime. However, small black holes can serve as catalysers and significantly speed up the phase transition. In fact, bubbles of true vacuum might already exist in our universe. We calculate the spectrum of Higgs particles produced by such a bubble, and use event generators to study their decay and subsequent evolution of the decay products to obtain the spectrum of emitted photons and neutrinos as a long-range signature. If the propagation of the bubble walls slows down due to interaction with the surrounding matter and plasma, these signals can reach us before the bubble wall hits us, thus representing the signals of the doomsday.
@@ 1 Introduction
It is believed that our universe has so far undergone a number of phase transitions at different energy scales, e.g. GUT, electroweak, QCD… Phase transitions usually introduce drastic changes in the structure of the universe, and if they happen late enough, they would almost certainly be fatal to any existing life forms. Until recently, late-time phase transitions have not been a topic of intensive study (see however [1, 2, 3]).
@@
After the Higgs discovery, it became apparent that a scenario similar to the one outlined in [2] (that the Higgs potential might develop instability with our current universe sitting in the unstable false vacuum) might actually be true [4]. This might not be an immediate problem, since the lifetime of our universe in the false vacuum could be many billions of years. However, as shown in [5, 6, 7, 8, 9, 10, 11], small primordial black holes can play the role of Higgs vacuum decay catalysers and significantly increase the tunneling probability. Depending on the small black hole mass and the exact parameters in the Higgs potential, the tunneling probability could be high enough to produce bubbles of true vacuum within the present lifetime of the universe.
@@
Once a bubble of true vacuum is created, it will expand at a speed close to the speed of light. A bubble wall sweeps through the universe and destroys (or modifies beyond recognition) everything it encounters. If the bubble wall is approaching with the speed of light, then no signal emitted from the wall can reach the observer before the wall, simply because the wall is already moving with the maximal possible speed. However, interaction with surrounding plasma and matter could slow down the wall propagation [12]. For example, if the bubble is created one million light years away from us (somewhere between us and Andromeda) and the wall slows down by only $1$ km/s, then the signal can reach us about three years before the bubble wall. It is then of utmost importance to calculate the characteristics of the possible signals that might come from an approaching bubble.
@@
We investigate here the process in which the Higgs tunnelling can leave a signature in the form of produced particles that can potentially reach us before the bubble. In particular, we calculate the spectrum of Higgs particles created due to the vacuum mismatch inside and outside of the bubble. Then, we use HDecay [13, 14] event generator to calculate the immediate Higgs decay products. Since we are mostly interested in obtaining the long range signature (i.e. emitted photons and neutrinos), we then use Pythia [15] to hadronize the decay products and obtain the final photon and neutrino energy spectra. To quantify the amount of particles that can reach us before the bubble wall, we introduce friction between the wall and the environment, which leads to a terminal velocity slower than the speed of light. We then integrate particle production up to the moment when the terminal velocity is reached and show that it is significant enough for a possible detection.
@@ 2 The Higgs vacuum decay
The standard model range of parameters is $0.01\lesssim\lambda_{*}\lesssim 0$ , $0.1M_{p}\lesssim\phi_{*}\lesssim M_{p}$, where, $M_{p}=\sqrt{\frac{1}{8\pi G}}=2.435\times 10^{18}$GeV, and $b\sim 10^{-4}$ [4, 9, 16]. This potential indicates that we live in an unstable vacuum, though with a very long lifetime [17].
As shown in [10] (for earlier work, see [2]), the inclusion of gravity can significantly modify the dynamics of the electroweak phase transition and drastically shorten the lifetime. We can parametrise the modification by including higher-order operators as
@@
Figure 1: This figure shows the Higgs potential $V(\phi)$ in Eq. (2.2). The true vacuum is at $\phi\approx 3.6\times 10^{-2}M_{p}$. The parameters are chosen to be within the standard model, i.e. $b=10^{-4}$, $\lambda=-0.001$, $\phi_{*}=0.5M_{p}$, and $\lambda_{6}=0.34$[16].
We adopt the conventions and follow the calculations in [10]. The nucleation rate in presence of gravity is determined by a bounce solution with Euclidean metric signature (+,+,+,+) with an action
$S_{E}=\int_{\mathcal{M}}\left[-\frac{1}{16\pi G}\mathcal{R}+\left(\frac{1}{2}g^{ab}\partial_{a}\phi\partial_{b}\phi+V(\phi)\right)\right]\sqrt{-g}d^{4}x.$ (2.3)
@@ 3 The true Higgs vacuum bubble propagation
As shown in [8, 9, 10], any primordial black hole lighter than $4\times 10^{14}$g at the time of formation would have evaporated by now, and in the absence of new physics beyond the standard model, would have entered the mass range in which it can trigger false vacuum decay (for related work see [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]). When the Higgs field tunnels through the vacuum barrier, a bubble of true vacuum is created and starts expanding. We assume that the geometry around the bubble is not far from a flat spacetime, so that we can neglect gravitational effects. In this case, the pure Higgs field , $\phi(\rho)$, action can be written as
@@
If we plot $r^{2}-t^{2}=\rho^{2}=$constant, we will find that the minimal $r$ happens at $t=0$. This is the moment when the bubble bounces back. Thus, the Higgs bubble is collapsing for $t<0$, bounces at $t=0$, and is expanding for $t>0$. The region $r>t$ is still in the false vacuum, while the $r<t$ region is in the true vacuum. So in this parametrization, nucleation happens at $t=0$, and then the true vacuum bubble expands. We simplify the problem by applying the thin wall approximation
@@
Creation of particles during the first-order phase transitions has been studied previously (e.g., [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]). In principle, particles can be produced via many different mechanisms. For our purpose, we will use the vacuum mismatch method [33, 39] which can be easily formulated in Minkowski (as opposed to Euclidean) space.
@@ 4 Particle production due to vacuum mismatch
The mismatch of vacua is usually accompanied by particle production. In our case, the false Higgs vacuum is outside the bubble, while the true vacuum is inside. As the bubble expands through space, the Higgs field vacuum abruptly changes from false to true. Bogolyiubov transformations then indicate that this vacuum mismatch will lead to the Higgs particle production. Since the Higgs particle is very heavy in the new vacuum, its decay products will be very energetic. For simplicity, we assume that the Higgs field tunnels to the true vacuum state homogeneously. This should be a good enough approximation since particles are produced locally, i.e. the relevant phase transition length scale is smaller than the size of the bubble at virtually any time after the initial bubble nucleation. Here, for the relevant length scale we take the width of the potential barrier in Fig. 1 ($\sim[10^{-2}M_{p}]^{-1}$), which is also the scale of the Higgs expectation value in the true vacuum.
@@
where $\tilde{\tau}^{*}$ is the characteristic time scale for the duration of the transition between the vacua. Since the number density of the produced particles crucially depends on this constant (see Eqs. (4.7) and (4.9)) we have to choose it in a meaningful way. We will set $\tilde{\tau}^{*}=-R_{0}$, where $R_{0}$ is the radius of the bubble at the time of nucleation. The justification for this choice is that this is the magnitude of the proper acceleration of the shell, i.e. $a=1/R_{0}$. The process of particle creation due to the vacuum mismatch is in this context closely related to Unruh radiation, which is determined by the magnitude of the proper acceleration. We note here that this magnitude is a Lorentz invariant quantity, $a=\sqrt{a_{\mu}a^{\mu}}$, which can be calculated from the equation of the bubble motion, $r^{2}-t^{2}=R_{0}^{2}$ (see Appendix-A [10]). As such, it is not equal to the coordinate acceleration that an outside observer would see. The proper acceleration is rather parametrized in terms of the proper time for an observer who is momentarily at rest with respect to the bubble wall. Since in our case $a=1/R_{0}$, wall expansion is a motion with a constant proper acceleration, and particles are produced continuously. Alternatively, we could take $\tilde{\tau}^{*}=-\eta^{-1}$, where $\eta\sim 10^{-2}M_{p}$ is the characteristic energy scale for the phase transition (i.e. the Higgs vacuum expectation value in the new vacuum, and also the width of the potential barrier). This choice will yield significantly more produced particles, however it lacks connection with the Unruh effect. We therefore go with the more conservative choice $\tilde{\tau}=-R_{0}$.
The solution for $h$ in Eq. (4.2) can be written as a combination of mode functions $g_{k}$, which satisfy $\triangledown^{2}g_{k}=-k^{2}$
@@
$N_{k}=\frac{B_{k}^{2}}{A_{k}^{2}-B_{k}^{2}}=\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}e^{4\omega_{+}R_{0}}-1,\right]^{-1}$ (4.9)
@@
Fig. 3 shows the number density of the produced Higgs particles per momentum mode per unit volume, $N_{k}=dN/(dVd^{3}\vec{k})$, as a function of their momenta, $k$, given in Planck units. The value of the Higgs’ mass in the false vacuum (where we currently live) is $M=125.09\,\text{GeV}$ [40]. For the choice of parameters we used in Fig. 1, the Higgs mass in the true vacuum is $\mu=7.16\times 10^{-4}M_{p}$. We emphasize that we do not have much freedom here, and this value might change by only one order of magnitude if we want to keep the values for $b,\lambda$, and $\phi_{*}$ within the standard model. For the same choice of parameters the bubble radius at the time of nucleation is $R_{0}=2\times 10^{4}M_{p}^{-1}$.
@@
$\mathcal{E}$ $=$ $\int_{0}^{\infty}\omega_{+}N_{k}\frac{d^{3}k}{(2\pi)^{3}}$ (4.10)
$=$ $\int_{0}^{\infty}\frac{\sqrt{\mu^{2}+k^{2}}}{\frac{(\sqrt{\mu^{2}+k^{2}}+\sqrt{M^{2}+k^{2}})^{2}}{(\sqrt{\mu^{2}+k^{2}}-\sqrt{M^{2}+k^{2}})^{2}}e^{4\sqrt{\mu^{2}+k^{2}}R_{0}}-1}$
$=$ $2.4\times 10^{-42}M_{p}^{4}=8.6\times 10^{31}\,\text{GeV}^{4}.$
@@ 5 Higgs decay into photons and neutrinos
A complete treatment of the Higgs decay is very complicated under these circumstances. Inside the bubble the Higgs field is in a new vacuum, so it is heavy, while outside is in the old vacuum, and it is light. A heavy Higgs cannot exist outside of the bubble, and it has to decay practically instantaneously. The decay products have to propagate outside in the old vacuum to reach a distant observer. Strictly speaking, one would have to model how the heavy Higgs propagates through the barrier and decays along the way. To simplify the task, we just assume that the Heavy Higgs decays instantaneously at the bubble wall. Ultimately, we want the spectrum of photons and neutrinos as a long-range observational signature. For that purpose, we will use the event generators that simulate the Higgs decay and subsequent evolution of the decay products. Generators like HDecay are convenient for our purpose since they evolve coupling constants with energy and can simulate heavy Higgs decay.
To simplify the analysis, we set $k=0$ in the definitions of $\omega_{+}$ and $\omega_{-}$, i.e. we consider the decay of Higgs particles at rest. Using this assumption, we calculate the number of Higgs particles produced due to vacuum mismatch from equation (4.9) for the specific value of the Higgs’s mass in true vacuum of $\mu=7.16\times 10^{-4}M_{p}$, resulting in $N_{k}(k=0)\simeq 1.365\times 10^{-25}$ (per Planck volume). Furthermore, we assume that the standard model (SM) remains valid up to the true vacuum range at $\phi\approx 3.6\times 10^{-2}M_{p}$. Our analysis includes only the leading-order contributions to Higgs decay to the fermions, disregarding higher-order corrections. This approximation is justified within the framework of the effective field theory, where the higher-order terms are highly suppressed at the mass scales under consideration.
To obtain the branching ratios and the total decay width of the SM Higgs particle at a true vacuum, we use HDECAY. We use only the tree level diagrams and switch off the electroweak corrections for the simulations because they become unstable beyond the 1 TeV range. For the rest of the input parameters, we follow the instructions given in [13] without any changes.
The output results for the branching ratios into fermions and bosons, as well as the total decay width, are listed in Tables 1 and 2, respectively.
Table 1: Branching ratios for the Higgs decays into fermions.
$M_{H}$ (GeV) $BR({b\bar{b}})$ $BR({c\bar{c}})$ $BR({s\bar{s}})$ $BR({t\bar{t}})$ $BR({\tau^{+}\tau^{-}})$ $BR({\mu^{+}\mu^{-}})$
$1.75\times 10^{15}$ $6.08\times 10^{-31}$ $2.85\times 10^{-32}$ $1.72\times 10^{-34}$ $1.57\times 10^{-27}$ $1.35\times 10^{-30}$ $4.77\times 10^{-33}$
Table 2: Branching ratios for the Higgs decays into bosons and the total decay width.
$M_{H}$ (GeV) $BR(GG)$ $BR(\gamma\gamma)$ $BR(Z\gamma)$ $BR(WW)$ $BR(ZZ)$ Width (GeV)
$1.75\times 10^{15}$ $1.08\times 10^{-04}$ $46\times 10^{-07} 9.9$$90\times 10^{-07}$ $6.95\times 10^{-01}$ $3.49\times 10^{-01}$ $2.30\times 10^{39}$
Note that the Higgs decay width can become larger than the Higgs mass, which indicates that the Higgs behaves more like a resonance than a well-defined particle at these scales, as argued in [41]. The output data consisting of the Higgs mass, branching ratios and the Higgs decay width, generated from HDecay, were used as input parameters into Pythia for hadronization (formation of mesons and baryons) and final-state decays. In Pythia, we successfully simulated hadronization and decay for various channels up to a Higgs mass of $1\times 10^{7}$ GeV without encountering simulation errors. We then extrapolated these findings (number of photons and neutrinos) using the logarithmic extrapolation method, to the true vacuum region at a Higgs mass of $7.16\times 10^{-4}M_{p}=1.75\times 10^{15}$ GeV, which provided the most accurate predictions when compared with theoretical decay width calculations for each decay channels. After extrapolation, we finally get the observational signature in terms of the final-state particles.
In the final state, the most abundant particles are photons and neutrinos, while stable particles such as electrons, muons, protons, neutrons, and other heavy baryons are produced in significantly smaller quantities. These final-state particles primarily originate from Higgs decays into quarks, which subsequently hadronize. Since photons and neutrinos vastly outnumber the stable particles, they represent the primary observational signature of Higgs vacuum decay, allowing us to neglect the contribution from stable particles. They anyway represent the desired long-range signature of the bubble formation.
Figures 4 and 5 show the spectra of photons and neutrinos respectively, i.e. their number densities as a function of their energies, originated in decay of $N_{k}(k=0)\simeq 1.365\times 10^{-25}$ Higgs particles per unit volume produced by a bubble of true vacuum.
Figure 4: The spectrum of photons created by a bubble of true vacuum due to the Higgs decay.
Figure 5: The spectrum of neutrinos created by a bubble of true vacuum due to the Higgs decay.
During the expansion of the bubble, vacuum is continuously changing from an old one to a new one, so particles are produced continuously in this process. Since the plots in 4 and 5 are actually particle number densities given in Planck units, to get the total number of produced particles (as a function of energy) we have to integrate over the volume of the bubble.
@@
We also note that the spectra in Figures 4 and 5 show the particle number density at the site of creation. To obtain the observed spectra on Earth, we have to propagate these particles through expanding space. The corresponding particle flux (number of particles per unit area per unit time) would therefore be diluted by a factor of $1/[4\pi d^{2}(1+z)]$, where $d$ is the physical distance between the source and the observer, while $z$ is the redshift of the source. The factor $(1+z)$ appears due to the relativistic time delay because particles hit the sphere with the radius $d$ less frequently (two particles emitted $\Delta t$ apart will be measured $(1+z)\Delta t$ apart. Accordingly, we have to correct the particle energies by a factor of $1+z$ because the individual particle’s energies are redshifted by that factor.

### v4

# The Signals of Doomsday I: False Higgs vacuum decay signatures

@@ Abstract
The measured standard model parameters indicate that we might live in a false Higgs vacuum, though with a very long lifetime. However, small black holes can serve as catalysers and significantly speed up the phase transition. In fact, bubbles of true vacuum might already exist in our universe. If the propagation of the bubble walls slows down due to interaction with the surrounding matter and plasma, these signals can reach us before the bubble wall hits us. Using the vacuum mismatch method, we calculate the spectrum of the Higgs particles produced by such a bubble until the terminal velocity is reached. In addition, we show that frictional dissipation at the terminal wall velocity generates a large population of thermally produced Higgs particles, which continues even after the mismatch channel shuts off. Since the Higgs is neutral, a good part of the final decay products (after hadronization, annihilation and decay of unstable particles) will be photons and neutrinos, which will then act as a long-range signature. For the conservative set of parameters used here, the thermal channel produces a macroscopically large burst of high–energy neutrinos and photons from Higgs decays, which could be detectable from sufficiently nearby bubbles with current or upcoming multi–messenger facilities.
@@ 1 Introduction
It is believed that our universe has so far undergone a number of phase transitions at different energy scales, e.g. GUT, electroweak, QCD… Phase transitions usually introduce drastic changes in the structure of the universe, and if they happen late enough, they would almost certainly be fatal to any existing life forms. Until recently, late-time phase transitions have not been a topic of intensive study (see however [101, 58, 68, 5, 50, 44, 45, 16, 10, 11, 97, 51]). A similar late–time analysis of $SU(3)_{c}$ symmetry breaking in a true–vacuum background has been presented in Ref. [96] and an analysis for $U(1)_{\rm EM}$ symmetry breaking will appear soon.
@@
After the Higgs discovery, it became apparent that a scenario similar to the one outlined in [58] (that the Higgs potential might develop instability with our current universe sitting in the unstable false vacuum) might actually be true [42, 48, 53, 99, 46]. This might not be an immediate problem, since the lifetime of our universe in the false vacuum could be many billions of years. However, as shown in [34, 33, 17, 60, 28, 29, 62, 30, 102], small primordial black holes can play the role of Higgs vacuum decay catalysers and significantly increase the tunneling probability. Depending on the small black hole mass and the exact parameters in the Higgs potential, the tunneling probability could be high enough to produce bubbles of true vacuum within the present lifetime of the universe.
@@
Once a bubble of true vacuum is created, it will expand at a speed close to the speed of light. A bubble wall sweeps through the universe and destroys (or modifies beyond recognition) everything it encounters. If the bubble wall is approaching with the speed of light, then no signal emitted from the wall can reach the observer before the wall, simply because the wall is already moving with the maximal possible speed. However, interaction with surrounding plasma and matter could slow down the wall propagation [37, 12, 75, 4]. For example, if the bubble is created one million light years away from us (somewhere between us and Andromeda) and the wall slows by only $1$ km/s, then the signal can reach us about three years before the bubble wall. It is therefore of utmost importance to calculate the characteristics of the possible signals that might come from an approaching bubble.
@@
We investigate here the process in which the Higgs tunneling can leave a signature in the form of produced particles that can potentially reach us before the bubble. In particular, we calculate the spectrum of the Higgs particles created due to the vacuum mismatch inside and outside of the bubble. If there is no friction with the environment, this mechanism will continue indefinitely during the bubble expansion and lead to enormous particle production. Since the bubble walls expand almost with the speed of light, the signal will hit us practically at the same time as the wall itself. Therefore, to quantify the amount of particles that can reach us before the bubble wall, we introduce friction between the wall and the environment, which leads to a terminal wall velocity slower than the speed of light. We then integrate particle production up to the moment when the terminal velocity is reached.
In addition to this vacuum–mismatch emission, the frictional dissipation at the terminal velocity produces a sustained thermal bath of Higgs (and other) particles in the shocked plasma behind the wall. This thermal channel remains active even after the direct mismatch mechanism shuts off, and can exceed the mismatch yield by many orders of magnitude. The Higgs particle is unstable, and it will quickly decay into the standard model particles. Since it is neutral, it will decay into an equal number of charged particles and antiparticles (i.e. quarks and antiquarks, bosons and antibosons, leptons and antileptons). After hadronization, decay of the unstable particles, and annihilation of particles and antiparticles, we will get mostly photons and neutrinos as the final decay products. It is thus safe to assume that most of the energy invested into the Higgs particle production will eventually end up as photons and neutrinos, which will represent a long range signature of the approaching bubble.
During the expansion of a bubble in a first-order phase transition, particles produced at the wall split into two kinematic populations: those that propagate outward into the false-vacuum exterior, and those that enter the true-vacuum interior. While particles falling inward may subsequently decay according to the true-vacuum symmetries, only a small subset of their light decay products can overtake the relativistically moving wall and re-emerge into the false vacuum. This contribution is therefore negligible. In the present work we focus exclusively on the outward-propagating population—namely, particles that decay in the vicinity of the bubble wall and whose decay products travel toward an observer in the false vacuum—consistently neglecting the suppressed inward-to-outward channel.
@@ 2 The Higgs vacuum decay
The standard model range of parameters is $0.01\lesssim\lambda_{*}\lesssim 0$ , $0.1M_{p}\lesssim\phi_{*}\lesssim M_{p}$, where, $M_{p}=\sqrt{\frac{1}{8\pi G}}=2.435\times 10^{18}\,\mathrm{GeV,}$ and $b\sim 10^{-4}$ [42, 28, 27]. This potential indicates that we live in an unstable vacuum, though with a very long lifetime [6].
As shown in [29] (for earlier work, see [58]), the inclusion of gravity can significantly modify the dynamics of the electroweak phase transition and drastically shorten the lifetime. We can parameterise the modification by including higher-order operators as
@@
Figure 1: This figure shows the Higgs potential $V(\phi)$ in Eq. (2.2). The true vacuum is at $\phi\approx 3.6\times 10^{-2}M_{p}$. The parameters are chosen to be within the standard model, i.e. $b=10^{-4}$, $\lambda_{*}=-0.001$, $\phi_{*}=0.5M_{p}$, and $\lambda_{6}=0.34$[27].
We adopt the conventions and follow the calculations in [29]. The nucleation rate in the presence of gravity is determined by a bounce solution with Euclidean metric signature (+,+,+,+) with an action
$S_{E}=\int_{\mathcal{M}}\left[-\frac{1}{16\pi G}\mathcal{R}+\left(\frac{1}{2}g^{ab}\partial_{a}\phi\partial_{b}\phi+V(\phi)\right)\right]\sqrt{g}d^{4}x.$ (2.3)
@@ 3 The true Higgs vacuum bubble propagation
As shown in [60, 28, 29], any primordial black hole lighter than $4\times 10^{14}$g at the time of formation would have evaporated by now, and in the absence of new physics beyond the standard model, would have entered the mass range in which it can trigger false vacuum decay (for related work see [36, 102, 32, 30, 24, 25, 23, 21, 22, 16, 50]). When the Higgs field tunnels through the vacuum barrier, a bubble of true vacuum is created and starts expanding. We assume that the geometry around the bubble is not far from a flat spacetime, so that we can neglect gravitational effects. In this case, the pure Higgs field, $\phi(\rho)$, action can be written as
@@
If we plot $r^{2}-t^{2}=\rho^{2}=$constant, we will find that the minimal $r$ happens at $t=0$. This is the moment when the bubble bounces back. Thus, the Higgs bubble is collapsing for $t<0$, bounces at $t=0$, and is expanding for $t>0$. The region $r>t$ is still in the false vacuum, while the $r<t$ region is in the true vacuum. So in this parametrization, nucleation happens at $t=0$, and then the true vacuum bubble expands. We simplify the problem by applying the thin wall approximation [^1] We model the wall as thin where $\ell_{\rm wall}$ denotes the physical thickness of the bubble wall, i.e. the characteristic distance over which the Higgs background interpolates from the false to the true vacuum. Parametrically this is set by the inverse mass of small fluctuations about the true vacuum, $\ell_{\rm wall}\sim[V^{\prime\prime}(\phi_{\rm true})]^{-1/2}\equiv 1/\mu$. For our benchmark $\mu=7.16\times 10^{-4}M_{p}$, so $\ell_{\rm wall}\simeq 1/\mu\simeq 1.40\times 10^{3}\,M_{p}^{-1}$. Comparing with the nucleation radius $R_{0}=2.0\times 10^{4}\,M_{p}^{-1}$ gives the thin-wall parameter $\varepsilon\equiv\ell_{\rm wall}/R_{0}\simeq(1.40\times 10^{3})/(2.0\times 10^{4})\simeq 0.07$, so curvature variation across the wall thickness is modest.
@@
Creation of particles during the first-order phase transitions has been studied previously (e.g., [112, 90, 89, 85, 86, 84, 105, 103, 63, 87, 72, 69, 52, 49, 98, 83]). In principle, particles can be produced via many different mechanisms. For our purpose, we will first use the vacuum mismatch method [84, 104] which can be easily formulated in Minkowski (as opposed to Euclidean) space, then we will use the thermal production mechanism in the later section.
@@ 4 Particle production due to vacuum mismatch
The mismatch of vacua is usually accompanied by particle production. In our case, the false Higgs vacuum is outside the bubble, while the true vacuum is inside. As the bubble expands through space, the Higgs field vacuum abruptly changes from false to true. Bogoliubov transformations then indicate that this vacuum mismatch will lead to the Higgs particle production. Since the Higgs particle is very heavy in the new vacuum, its decay products will be very energetic. For simplicity, we assume that the Higgs field tunnels to the true vacuum state homogeneously. This should be a good enough approximation since particles are produced locally, i.e. the relevant phase transition length scale is smaller than the size of the bubble at virtually any time after the initial bubble nucleation. Here, for the relevant length scale we take the width of the potential barrier in Fig. 1 ($\sim[10^{-2}M_{p}]^{-1}$), which is also the scale of the Higgs expectation value in the true vacuum.
@@
where $\tilde{\tau}^{*}$ is the characteristic time scale for the duration of the transition between the vacua. Since the number density of the produced particles crucially depends on this constant (see Eqs. (4.7) and (4.9)) we have to choose it in a meaningful way. We will set $\tilde{\tau}^{*}=-R_{0}$, where $R_{0}$ is the radius of the bubble at the time of nucleation. The justification for this choice is that this is the magnitude of the proper acceleration of the shell, i.e. $a=1/R_{0}$. The process of particle creation due to the vacuum mismatch is in this context closely related to Unruh radiation, which is determined by the magnitude of the proper acceleration. We note here that this magnitude is a Lorentz invariant quantity, $a=\sqrt{a_{\mu}a^{\mu}}$, which can be calculated from the equation of the bubble motion, $r^{2}-t^{2}=R_{0}^{2}$ (see Appendix-A [A]). As such, it is not equal to the coordinate acceleration that an outside observer would see. The proper acceleration is rather parametrized in terms of the proper time for an observer who is momentarily at rest with respect to the bubble wall. Since in our case $a=1/R_{0}$, wall expansion is a motion with a constant proper acceleration, and particles are produced continuously. Alternatively, we could take $\tilde{\tau}^{*}=-\kappa^{-1}$, where $\kappa\sim 10^{-2}M_{p}$ is the characteristic energy scale for the phase transition (i.e. the Higgs vacuum expectation value in the new vacuum, and also the width of the potential barrier). This choice will yield significantly more produced particles, however it lacks connection with the Unruh effect. We therefore go with the more conservative choice $\tilde{\tau}^{*}=-R_{0}$.
The solution for $h$ in Eq. (4.2) can be written as a combination of mode functions $g_{k}$, which satisfy $\nabla^{2}g_{k}=-k^{2}$
@@
$N_{k}=\frac{B_{k}^{2}}{A_{k}^{2}-B_{k}^{2}}=\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}e^{4\omega_{+}R_{0}}-1\right]^{-1}$ (4.9)
@@
Fig. 3 shows the number density of the produced Higgs particles per momentum mode per unit volume, $N_{k}=dN/(dVd^{3}\vec{k})$, as a function of their momenta, $k$, given in Planck units. The value of the Higgs’ mass in the false vacuum (where we currently live) is $M=125.09\,\mathrm{GeV}$ [1]. For the choice of parameters we used in Fig. 1, the Higgs mass in the true vacuum is $\mu=7.16\times 10^{-4}M_{p}$. We emphasize that we do not have much freedom here, and this value might change by only one order of magnitude if we want to keep the values for $b,\lambda$, and $\phi_{*}$ within the standard model. For the same choice of parameters the bubble radius at the time of nucleation is $R_{0}=2\times 10^{4}M_{p}^{-1}$.
@@
$\mathcal{E}$ $=$ $\int_{0}^{\infty}\omega_{+}N_{k}\frac{d^{3}\vec{k}}{(2\pi)^{3}}$ (4.10)
$=$ $\int_{0}^{\infty}\frac{\sqrt{\mu^{2}+k^{2}}}{\frac{(\sqrt{\mu^{2}+k^{2}}+\sqrt{M^{2}+k^{2}})^{2}}{(\sqrt{\mu^{2}+k^{2}}-\sqrt{M^{2}+k^{2}})^{2}}e^{4\sqrt{\mu^{2}+k^{2}}R_{0}}-1}\frac{d^{3}\vec{k}}{(2\pi)^{3}}$
$=$ $2.4\times 10^{-42}M_{p}^{4}=8.6\times 10^{31}\,\mathrm{GeV}^{4}.$
@@
A complete treatment of the Higgs decay is very complicated under these circumstances. Inside the bubble the Higgs field is in a new vacuum, so it is heavy, while outside is in the old vacuum, and it is light. A heavy Higgs cannot exist outside of the bubble, and it has to decay practically instantaneously. The decay products have to propagate outside in the old vacuum to reach a distant observer. Strictly speaking, one would have to model how the heavy Higgs propagates through the barrier and decays along the way. To simplify the task, we just assume that the heavy Higgs decays instantaneously at the bubble wall.
@@
During the expansion of the bubble, vacuum is continuously changing from an old one to a new one, so particles are produced continuously in this process. Since the plot in Figure 3 is actually particle number density given in Planck units, to get the total number of produced particles (as a function of energy) we have to integrate over the volume of the bubble. We also note that the spectra in Figure 3 show the particle number density at the site of creation. To obtain the observed spectra on Earth, we have to propagate these particles through expanding space. The corresponding particle flux (number of particles per unit area per unit time) would therefore be diluted by a factor of $1/[4\pi d^{2}(1+z)]$, where $d$ is the physical distance between the source and the observer, while $z$ is the redshift of the source. The factor $(1+z)$ appears due to the relativistic time delay because particles hit the sphere with the radius $d$ less frequently (two particles emitted $\Delta t$ apart will be measured $(1+z)\Delta t$ apart. Accordingly, we have to divide the particle energies by a factor of $1+z$ because the individual particle’s energies are redshifted by that factor.